# Regist3R: Incremental Registration with Stereo Foundation Model

### v1

@@ 3.4 Inference Strategy
When inference, previous approaches exhaustively perform pairwise inference on all image pairs [12] or a subset of pairs [16], resulting in computational bottleneck. Instead, we follow incremental registration paradigm [5], requiring only one inference for each new view, thus being highly efficient even when the scene scales larger. Specifically, we build a minimum spanning tree (MST) based on the view similarity, and perform pairwise inference only between the parents and children. This takes only $N-1$ times inference for a collection containing $N$ views. As the drifting happens when the tree goes deep, we propose a tree compression trick to reduce the length of reconstruction chain.
3.4.1 Incremental Registration
@@
For each registration step, traditional approaches, etc. COLMAP [5], select the next images based on the matching with registered images. While matching is not explicitly defined in Regist3R, we resort to image retrieval to select the next image. Following MASt3R-SfM [16], we use Aggregated Selective Matching Kernel (ASMK) [45] for efficient pairwise similarity measurement, which has shown excellent performance for retrieval. ASMK receives the feature map from the MASt3R [15]’s image encoder and outputs a similarity matrix. Then we build a minimum spanning tree (MST) and select the image with the largest summation of similarity with other images as the root [20]. As the pointmap of the root is required for bootstrapping, we adopt DUSt3R and select the root and its first child as the initial pair for pointmap estimation. Other approaches, like monocular MoGe [46] or multi-view VGGT [19], are also applicable. Then we set camera coordinate system of root view as the global coordinate system, and travel the MST to estimate the pointmap of each view.
3.4.2 Tree Depth Compression
@@
MST keeps the edge set with minimal cost, when the images are collected in a sequence, the MST tends to degenerate into a chain, which causes the drifting. Light3R-SfM [20] proposes to replace MST with a shortest path tree (SPT) to reduce the depth, but it leads to a large disparity between pairs when the scene views are sparse. Inspired by the key frame selection in Spann3R [22], we extend it to tree structure as key layer selection. Specifically, assuming that the key layer selection interval is $K$, the descendant nodes from the $nK+1^{th}$ layer to the $(n+1)K^{th}$ layer select their ancestors at the $nK^{th}$ layer as the reference view. For implementation, we change the parent node of the even-layer node to the grandparent node to achieve $K=2$ compression. Repeating it achieves power-of-2 compression. For dense view, we can set aggressive compression while for sparse view, we can set it smaller.
3.4.3 Pose Estimation
@@
Figure 3: Qualitative comparison of different methods across multiple scenes of CS-Drone3D dataset.
4 Experiments
@@
Training Data. Following DUSt3R, we train our model on a mix of 7 datasets: MegaDepth [47], ARKitScenes [48], Static Scenes 3D [49], Blended MVS [50], ScanNet++ [51], CO3D-v2 [52], and Waymo [53], covering the indoor, outdoor, and object-centric scenes. Training details can be found in Appendix Sec. 1.
Baselines. The relative baselines can be categorized into post-optimization-based and inference-only. For post-optimization-based approaches, DUSt3R [12] uses pointmap global alignment to align pairs. MASt3R-SfM [16] uses parse matching points alignment to improve scalability. This kind of approach achieves high accuracy pose estimation but takes a long time. For inference-only approaches, which are our main focus, Spann3R [22] organizes the views into a sequence and uses a memory bank to store history information. It requires the input views to be ordered. For an unordered set, the dense pairwise graph is built with DUSt3R, and a sequence is extracted according to the confidence. Fast3R [17] is a recently proposed multi-view extension of DUSt3R, which receives all views as input and predicts the pointmap of each view. MV-DUSt3R [18] is similar to Fast3R but Fast3R achieves better performance, therefore we only take Fast3R as the representative of multi-view models.
We further refer to the infer-then-align mode of Light3R-SfM [20] to optimize the multi-view reconstruction efficiency and scalability of DUSt3R, which uses the common view to perform cross-pair Procrustes alignment after performing pairwise inference. We name this model DUSt3R.
Evaluation Datasets and Metrics. We evaluate our model on three public datasets: DTU [54], NRGBD [55], and 7scenes [56], and a self-made dataset CS-Drone3D. DTU, NRGBD, and 7scenes are indoor or object-centric datasets with ground truth depth maps and camera poses. CS-Drone3D is an oblique aerial dataset with a large spatial span, acquired by a DJI drone at an altitude of about 140 meters. It contains three scenes: field with 258 images, hotel with 197 images, and bridge with 275 images. Some examples are shown in Appendix Sec. 2.
Given a set of images, we follow previous works to compute the relative camera pose errors for all image pairs and measure the percentage of pairs with angular rotation/translation error below a certain threshold $\tau$, denoted as relative rotation accuracy (RRA@$\tau$) and relative translation accuracy (RTA@$\tau$). We also report mean Average Accuracy below 30 (mAA@30) defined as the area under the curve in (RRA@$\tau$, RTA@$\tau$) at a threshold $\tau$, integrated over [1, 30]. We report its accuracy score average over all data samples. We also report the runtime on a system with an NVIDIA A100 (40GB). For the datasets that provide depth maps, we further report the point cloud precision with the measurement of point cloud accuracy (Acc) and completion (Comp). We shuffle the images to evaluate the performance on unordered set.
4.1 Comparative Study
4.1.1 Object-level reconstruction.
@@ 4.1 Comparative Study
Table 1: Performance comparison on DTU dataset
@@
Table 2: Performance comparison on Neural RGBD dataset
@@
Table 3: Performance comparison on 7 scenes dataset
@@
Table 4: Performance Comparison on CS-Drone3D Dataset
@@
4.1.2 Indroor scenes reconstruction.
@@
4.1.3 Outdoor aerial reconstruction.
@@
Figure 4: Performance of Regist3R with different tree compression factors. The depth of the tree is compressed to $1/2^{K}$ of the original depth.
4.2 Ablation Study
@@
Table 5: Ablation Study on confidence-aware autoregressive training. The w/o AR is evaluated with the same model but the confidence values are filled with all ones.
@@
Table 6: Ablation study on tree type
@@
4.2.1 Confidence-aware autoregressive training
@@
4.2.2 Tree building strategy
Previous work [20] proposes to use the shortest path tree (SPT) instead of the commonly used minimum spanning tree (MST). However, our experiments show that this method fails to reconstruct scenes due to large spacing between pairs in scenarios with large spatial spans. The results are shown in Tab. 6, where the tree compression factor is set to 1 for MST. The MST achieves better performance than SPT on all metrics. The MST is more robust to the large spacing between pairs and is more suitable for outdoor aerial reconstruction.
@@ 5 Conclusion
There remains substantial work to be done towards modern incremental reconstruction. For example, the pointmap representation should be extended to support the derivation of intrinsics, extrinsics, and depth, removing the constraint of intrinsic sharing. The heuristic tree-building strategy should be replaced by a more optimal method. Modern bundle adjustment should be developed to address the drifting that inherently exists in incremental registration. The combination of the global multi-view model and incremental stereo model should be explored to achieve truly scalable and accurate multi-view reconstruction. We hope our work can inspire more research in this direction.
@@ Appendix A Training Details
In our experiment, we use two ViT-L [57] transformers for the reference encoder and target encoder, and two ViT-B transformers for the reference decoder and target decoder, as well as a DPT output head for target view pointmap regression. As we freeze the target encoder, the modules actually trained are one ViT-L encoder, two ViT-B decoders, and a DPT head.
@@ Appendix B Description of CS-Drone3D Dataset
CS-Drone3D is an oblique aerial dataset with a large spatial span, acquired by a DJI drone at an altitude of about 140 meters. It contains three scenes: field with 258 images, hotel with 197 images, and bridge with 275 images. The camera pose and intrinsic parameters are estimated by commercial software, and the images are undistorted as a pinhole camera model. The resolution is downsampled to 960x716 for easy distribution.
@@ Appendix C Post-optimization for Regist3R
$\min_{\begin{subarray}{c}\{R_{i},t_{i}\}\\ \{R^{\prime}_{k},t^{\prime}_{k},s_{k}\}\end{subarray}}\sum_{k=1}^{K}\sum_{i=1}^{N}w_{k,i}\left[d_{R}(R_{i},R^{\prime}_{k}R_{k,i})+\left\|t_{i}-(s_{k}R^{\prime}_{k}t_{k,i}+t^{\prime}_{k})\right\|^{2}\right],$ (9)
@@
$w(d)=\exp\left(\frac{-5d}{\max_{k,i}(d_{k,i})}\right).$ (10)
@@
$d_{R}(R_{1},R_{2})=\left\|\log(R_{1}^{T}R_{2})\right\|^{2}.$ (11)
The roots of spanning tree need to be distinct. In practice, we adopt K-medoids [58] on similarity matrix to select the roots.
We evaluate the performance on DTU [54], NRGBD [55], 7 scenes [56] and CS-Drone3D, the results are reported in Tab. 7. We ensemble $K=3$ sequences for all datasets. In most of scenes, the pose precision consistently improved after ensemble, showing the effectiveness. But hotel scene in CS-Drone3D dataset is an exception, as its inherent complexity leads to significant variations in reconstruction results when selecting different root nodes for initialization. This instability ultimately causes the ensemble-based optimization to fail.
Table 7: Ablation on post-optimization. The results are evaluated on DTU, NRGBD, 7 Scenes and CS-Drone3D dataset. Regist3R with post-optimization is denoted with Regist3R* in the table. We ensemble $K=3$ sequences for all datasets.
@@
Figure 5: Illustration of the reconstruction results of Rubble scene from Hill19 dataset [59].
@@ Appendix D Reconstruction from 1000+ Views
Scene reconstruction and camera pose estimation from large image collections is a long standing challenge. Traditional approaches like COLMAP [5] often takes hours to days for reconstruction of collections containing thousands of views. Recent pointmap-based approaches achieves efficient reconstruction via model inference. But they still suffer from costly global alignment [12, 16], model generalizability or high memory consumption [18, 17, 19]. Therefore previous pointmap-based approaches fail to reconstruct a scene with 1000+ views. In this section, we demonstrate that Regist3R achieved this in a few minutes, taking Rubble from Hill19 dataset [59] as an example, which contains 1678 images, covering an area of about 1km2.
@@
Specifically, we uniformly sample key frames at equidistant intervals and generate pointmaps for each with pretrained VGGT [19] model. These key frames constitute the root node set. During forest construction, we first introduce a virtual node with zero-distance connections to all root nodes. Crucially, we assign infinite distances to edges between root nodes to prevent their inclusion in the MST. Then an MST is constructed starting from the virtual node, which is subsequently removed to obtain the final MST forest. Finally, we perform Regist3R inference based on these spanning trees to register the remaining views. The entire reconstruction process is completed in approximately 5 minutes on one A100 GPU. An illustration of the reconstruction is shown in Fig. 5.
@@
Figure 6: Number of keyframes versus camera pose accuracy.
Table 8: Total Time, Net Time, and FPS Analysis
@@ Appendix E Confidence Visualization
Figure 7: Visualization of the confidence decay. The first row shows the input images, the following two rows are the confidence heatmaps when the confidence-aware autoregression is enabled, with the first row showing the forward reconstruction sequence (from left to right) and the second row showing the backward (from right to left). The last two rows is the confidence heatmaps when the confidence-aware autoregression is disabled (input confidence is set to one). Obvious confidence decay can be observed when the autoregression is enabled, while the confidence is uniformly distributed when the autoregression is disabled.
Table 9: Performance Comparison of Confidence Strategies
@@
As shown in the ablation study, the confidence-aware autoregressive training improves the performance of Regist3R. This section quantitatively analyze it by visualizing the confidences, taking 10 frames from a video captured by hand-holded camera from LocalRF dataset [60] for evaluation, as shown in Fig. 7.

### v2

@@ 3.4. Inference Strategy
When inference, previous approaches exhaustively perform pairwise inference on all image pairs (Wang et al., 2024b) or a subset of pairs (Duisterhof et al., 2024), resulting in computational bottleneck. Instead, we follow incremental registration paradigm (Schonberger and Frahm, 2016), requiring only one inference for each new view, thus being highly efficient even when the scene scales larger. Specifically, we build a minimum spanning tree (MST) based on the view similarity, and perform pairwise inference only between the parents and children. This takes only N1N-1 times inference for a collection containing $N$ views. As the drifting happens when the tree goes deep, we propose a tree compression trick to reduce the length of reconstruction chain.
3.4.1. Incremental Registration
@@
For each registration step, traditional approaches, etc. COLMAP (Schonberger and Frahm, 2016), select the next images based on the matching with registered images. While matching is not explicitly defined in Regist3R, we resort to image retrieval to select the next image. Following MASt3R-SfM (Duisterhof et al., 2024), we use Aggregated Selective Matching Kernel (ASMK) (Tolias et al., 2013) for efficient pairwise similarity measurement, which has shown excellent performance for retrieval. ASMK receives the feature map from the MASt3R (Leroy et al., 2024)’s image encoder and outputs a similarity matrix. Then we build a minimum spanning tree (MST) and select the image with the largest summation of similarity with other images as the root (Elflein et al., 2025). As the pointmap of the root is required for bootstrapping, we adopt DUSt3R and select the root and its first child as the initial pair for pointmap estimation. Other approaches, like monocular MoGe (Wang et al., 2024c) or multi-view VGGT (Wang et al., 2025), are also applicable. Then we set camera coordinate system of root view as the global coordinate system, and travel the MST to estimate the pointmap of each view.
3.4.2. Tree Depth Compression
@@
MST keeps the edge set with minimal cost, when the images are collected in a sequence, the MST tends to degenerate into a chain, which causes the drifting. Light3R-SfM (Elflein et al., 2025) proposes to replace MST with a shortest path tree (SPT) to reduce the depth, but it leads to a large disparity between pairs when the scene views are sparse. Inspired by the key frame selection in Spann3R (Wang and Agapito, 2024), we extend it to tree structure as key layer selection. Specifically, assuming that the key layer selection interval is $K$, the descendant nodes from the $nK+1^{th}$ layer to the $(n+1)K^{th}$ layer select their ancestors at the $nK^{th}$ layer as the reference view. For implementation, we change the parent node of the even-layer node to the grandparent node to achieve $K=2$ compression. Repeating it achieves power-of-2 compression. For dense view, we can set aggressive compression while for sparse view, we can set it smaller.
3.4.3. Pose Estimation
@@
Figure 4. Qualitative comparison of different methods across multiple scenes of CS-Drone3D dataset.
4. Experiments
@@
Training Data. Following DUSt3R, we train our model on a mix of 7 datasets: MegaDepth (Li and Snavely, 2018), ARKitScenes (Baruch et al., 2021), Static Scenes 3D (Schröppel et al., 2022), Blended MVS (Yao et al., 2020), ScanNet++ (Yeshwanth et al., 2023), CO3D-v2 (Reizenstein et al., 2021), and Waymo (Sun et al., 2020), covering the indoor, outdoor, and object-centric scenes. Training details can be found in Appendix Sec. 1.
Baselines. The relative baselines can be categorized into post-optimization-based and inference-only. For post-optimization-based approaches, DUSt3R (Wang et al., 2024b) uses pointmap global alignment to align pairs. MASt3R-SfM (Duisterhof et al., 2024) uses parse matching points alignment to improve scalability. This kind of approach achieves high accuracy pose estimation but takes a long time. For inference-only approaches, which are our main focus, Spann3R (Wang and Agapito, 2024) organizes the views into a sequence and uses a memory bank to store history information. It requires the input views to be ordered. For an unordered set, the dense pairwise graph is built with DUSt3R, and a sequence is extracted according to the confidence. Fast3R (Yang et al., 2025) is a recently proposed multi-view extension of DUSt3R, which receives all views as input and predicts the pointmap of each view. MV-DUSt3R (Tang et al., 2024) is similar to Fast3R but Fast3R achieves better performance, therefore we only take Fast3R as the representative of multi-view models.
We further refer to the infer-then-align mode of Light3R-SfM (Elflein et al., 2025) to optimize the multi-view reconstruction efficiency and scalability of DUSt3R, which uses the common view to perform cross-pair Procrustes alignment after performing pairwise inference. We name this model DUSt3R.
Evaluation Datasets and Metrics. We evaluate our model on three public datasets: DTU (Aanæs et al., 2016), NRGBD (Azinović et al., 2022), and 7scenes (Shotton et al., 2013), and a self-made dataset CS-Drone3D. DTU, NRGBD, and 7scenes are indoor or object-centric datasets with ground truth depth maps and camera poses. CS-Drone3D is an oblique aerial dataset with a large spatial span, acquired by a DJI drone at an altitude of about 140 meters. It contains three scenes: field with 258 images, hotel with 197 images, and bridge with 275 images. Some examples are shown in Appendix Sec. 2.
[…]
4.1. Comparative Study
4.1.1. Object-level reconstruction.
@@ 4.1. Comparative Study
Table 1. Performance comparison on DTU dataset
@@
Table 2. Performance comparison on Neural RGBD dataset
@@
Table 3. Performance comparison on 7 scenes dataset
@@
[…]
Table 4. Performance Comparison on CS-Drone3D Dataset
@@
[…]
4.1.2. Indroor scenes reconstruction.
@@
4.1.3. Outdoor aerial reconstruction.
@@
Figure 5. Performance of Regist3R with different tree compression factors. The depth of the tree is compressed to $1/2^{K}$ of the original depth.
4.2. Ablation Study
@@
Table 5. Ablation Study on confidence-aware autoregressive training. The w/o AR is evaluated with the same model but the confidence values are filled with all ones.
@@
Table 6. Ablation study on tree type
@@
4.2.1. Confidence-aware autoregressive training
@@
4.2.2. Tree building strategy
Previous work (Elflein et al., 2025) proposes to use the shortest path tree (SPT) instead of the commonly used minimum spanning tree (MST). However, our experiments show that this method fails to reconstruct scenes due to large spacing between pairs in scenarios with large spatial spans. The results are shown in Tab. 6, where the tree compression factor is set to 1 for MST. The MST achieves better performance than SPT on all metrics. The MST is more robust to the large spacing between pairs and is more suitable for outdoor aerial reconstruction.
@@ 5. Conclusion
There remains substantial work to be done towards modern incremental reconstruction. For example, the heuristic tree-building strategy should be replaced by a more optimal method. Modern bundle adjustment should be developed to address the drifting that inherently exists in incremental registration. The combination of the global multi-view model and incremental stereo model should be explored to achieve truly scalable and accurate multi-view reconstruction. We hope our work can inspire more research in this direction.
Acknowledgements.
This work is supported by the National Nature Science Foundation of China (NSFC) under Grant No. 61902415, and the Open Fund of Science and Technology on Parallel and Distributed Processing Laboratory (PDL) under Grant No. WDZC20235250106.
@@ Appendix A Training Details
In our experiment, we use two ViT-L (Ranftl et al., 2021) transformers for the reference encoder and target encoder, and two ViT-B transformers for the reference decoder and target decoder, as well as a DPT output head for target view pointmap regression. As we freeze the target encoder, the modules actually trained are one ViT-L encoder, two ViT-B decoders, and a DPT head.
@@
Figure 6. Example images from CS-Drone3D dataset. The first row shows the field scene, the second row shows the hotel scene, and the last row shows the bridge scene. The images are acquired by a DJI drone at an altitude of about 140 meters.
@@ Appendix B Description of CS-Drone3D Dataset
CS-Drone3D is an oblique aerial dataset with a large spatial span, acquired by a DJI drone at an altitude of about 140 meters. It contains three scenes: field with 258 images, hotel with 197 images, and bridge with 275 images. The camera pose and intrinsic parameters are estimated by commercial software, and the images are undistorted as a pinhole camera model. The resolution is downsampled to 960x716 for easy distribution. Some examples are shown in Fig. 6.
@@ Appendix C Post-optimization for Regist3R
(9) $\min_{\begin{subarray}{c}\{R_{i},t_{i}\}\\ \{R^{\prime}_{k},t^{\prime}_{k},s_{k}\}\end{subarray}}\sum_{k=1}^{K}\sum_{i=1}^{N}w_{k,i}\left[d_{R}(R_{i},R^{\prime}_{k}R_{k,i})+\left\|t_{i}-(s_{k}R^{\prime}_{k}t_{k,i}+t^{\prime}_{k})\right\|^{2}\right],$
@@
(10) $w(d)=\exp\left(\frac{-5d}{\max_{k,i}(d_{k,i})}\right).$
@@
(11) $d_{R}(R_{1},R_{2})=\left\|\log(R_{1}^{T}R_{2})\right\|^{2}.$
The roots of spanning tree need to be distinct. In practice, we adopt K-medoids (Schubert and Rousseeuw, 2019) on similarity matrix to select the roots.
We evaluate the performance on DTU (Aanæs et al., 2016), NRGBD (Azinović et al., 2022), 7 scenes (Shotton et al., 2013) and CS-Drone3D, the results are reported in Tab. 7. We ensemble $K=3$ sequences for all datasets. In most of scenes, the pose precision consistently improved after ensemble, showing the effectiveness. But hotel scene in CS-Drone3D dataset is an exception, as its inherent complexity leads to significant variations in reconstruction results when selecting different root nodes for initialization. This instability ultimately causes the ensemble-based optimization to fail.
Table 7. Ablation on post-optimization. The results are evaluated on DTU, NRGBD, 7 Scenes and CS-Drone3D dataset. Regist3R with post-optimization is denoted with Regist3R* in the table. We ensemble $K=3$ sequences for all datasets.
@@
Figure 7. Illustration of the reconstruction results of Rubble scene from Hill19 dataset (Turki et al., 2022).
@@ Appendix D Reconstruction from 1000+ Views
Scene reconstruction and camera pose estimation from large image collections is a long standing challenge. Traditional approaches like COLMAP (Schonberger and Frahm, 2016) often takes hours to days for reconstruction of collections containing thousands of views. Recent pointmap-based approaches achieves efficient reconstruction via model inference. But they still suffer from costly global alignment (Wang et al., 2024b; Duisterhof et al., 2024), model generalizability or high memory consumption (Tang et al., 2024; Yang et al., 2025; Wang et al., 2025). Therefore previous pointmap-based approaches fail to reconstruct a scene with 1000+ views. In this section, we demonstrate that Regist3R achieved this in a few minutes, taking Rubble from Hill19 dataset (Turki et al., 2022) as an example, which contains 1678 images, covering an area of about 1km2.
@@
Specifically, we uniformly sample key frames at equidistant intervals and generate pointmaps for each with pretrained VGGT (Wang et al., 2025) model. These key frames constitute the root node set. During forest construction, we first introduce a virtual node with zero-distance connections to all root nodes. Crucially, we assign infinite distances to edges between root nodes to prevent their inclusion in the MST. Then an MST is constructed starting from the virtual node, which is subsequently removed to obtain the final MST forest. Finally, we perform Regist3R inference based on these spanning trees to register the remaining views. The entire reconstruction process is completed in approximately 5 minutes on one A100 GPU. An illustration of the reconstruction is shown in Fig. 7.
@@
Figure 8. Number of keyframes versus camera pose accuracy.
Table 8. Total Time, Net Time, and FPS Analysis
@@ Appendix E Confidence Visualization
Figure 9. Visualization of the confidence decay. The first row shows the input images, the following two rows are the confidence heatmaps when the confidence-aware autoregression is enabled, with the first row showing the forward reconstruction sequence (from left to right) and the second row showing the backward (from right to left). The last two rows is the confidence heatmaps when the confidence-aware autoregression is disabled (input confidence is set to one). Obvious confidence decay can be observed when the autoregression is enabled, while the confidence is uniformly distributed when the autoregression is disabled.
Table 9. Performance Comparison of Confidence Strategies
@@
As shown in the ablation study, the confidence-aware autoregressive training improves the performance of Regist3R. This section quantitatively analyze it by visualizing the confidences, taking 10 frames from a video captured by hand-holded camera from LocalRF dataset (Meuleman et al., 2023) for evaluation, as shown in Fig. 9.